OpenStudy (anonymous):

Solve by substitution 15x-20y=1 15y=-2+5x

OpenStudy (mertsj):

\[y=\frac{-2+5x}{15}\]

OpenStudy (mertsj):

\[15x-20(\frac{-2+5x}{15})=1\] \[15x-4(\frac{-2+5x}{3})=1\] \[15x+\frac{8-20x}{3}=1\] \[45x+8-20x=3\] \[25x=-5\] \[x=\frac{-1}{5}\]

OpenStudy (mertsj):

\[y=\frac{-2+5x}{15}\] \[y=\frac{-2+5(\frac{-1}{5})}{15}=\frac{-2-1}{15}=\frac{-3}{15}=\frac{-1}{5}\]

OpenStudy (mertsj):

\[(\frac{-1}{5},\frac{-1}{5})\]